# TokenSelect: Efficient Long-Context Inference and Length Extrapolation for LLMs via Dynamic Token-Level KV Cache Selection

### v2

@@ 6.6 Scaling Beyond 1 Million Context Length
7 Conclusion
In this paper, we introduce TokenSelect, a training-free approach for efficient long-context inference and length extrapolation. TokenSelect addresses the two major challenges faced by LLMs in processing long texts: the context length limitation from pre-training and the computational complexity of attention. This is achieved through a novel token-level selective sparse attention mechanism. Experimental results demonstrate that TokenSelect can achieve up to 23.84×23.84$\times$23.84 × speedup in attention computation and up to 2.28×2.28$\times$2.28 × acceleration in end-to-end inference latency, while exhibiting superior performance across multiple long-context benchmarks.
8 Limitations
Our approach has inherent limitations that present opportunities for future work. A primary limitation of our method is that its training-free design—a significant advantage—acts as a double-edged sword, as its absolute performance is inherently tied to the quality of the underlying LLMs. Although our experiments demonstrate robustness of TokenSelect across various LLMs, some inherent shortcomings—such as the misrecognition of UUID strings by Yi-1.5-6B-Chat—indicate that certain issues may still require training to resolve. Moreover, while our method currently achieves state-of-the-art performance in long-context inference, recent long-text post-training techniques in the LLM community have shown impressive performance; notably, our TokenSelect is orthogonal to these approaches and can be employed during inference to trade a slight performance drop for significant efficiency gains. Finally, although our method achieves state-of-the-art efficiency improvements in long-context inference, the task remains inherently resource-intensive. For instance, even with an 8B-parameter model, complex benchmarks (e.g., RULER) can require approximately 8$\times$×A100 GPUs for nearly one day of runtime, and the computational cost is expected to increase substantially for larger models. We hope that our work, together with the community’s advances in model design, algorithm development, and infrastructure optimization, will help pave the way for further mitigating these computational challenges.
Appendix A Formal Statement and Proof of Lemma
@@
$\min_{1\leq i\leq k}{\hat{\mathbf{q}}_{2}\mathbf{k}_{i}}$ $-\max_{j>k}{\hat{\mathbf{q}}_{2}\mathbf{k}_{j}}\geq\hat{\mathbf{q}}_{1}\mathbf{k}_{p}\cos{\theta}-\|\mathbf{k}\|_{\text{max}}\sin{\theta}$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ -(\hat{\mathbf{q}}_{1}\mathbf{k}_{p+1}\cos{\theta}+\|\mathbf{k}\|_{\text{max}}\sin{\theta})$
@@
In order to have Eqn. (11) $>0$, we require
@@
[…]
@@
Appendix B Detailed Descriptions on Baselines
@@
  • StreamingLLM Xiao et al. (2024b): The state-of-the-ar method for long-context inference with predefined sparse patterns. Similar approaches include LM-Infinite Han et al. (2024).
@@
  • MInference Jiang et al. (2024): The state-of-the-ar method for long-context prefilling acceleration, utilizing three sparse patterns including block-level sparse attention.
Appendix C More Information on Datasets
@@
Appendix D Effiency Comparison with MInference
@@
Length FlashAttention-2 (vLLM) MInference (vLLM) TokenSelect
@@
Appendix E Experimental Results on LongBench

### v4

@@ B.1 Scaling Beyond 1 Million Context Length
B.2 Scaling to 72 Billion Parameters
To demonstrate the scalability of our approach to larger models, we conducted additional experiments using Qwen2-72B-Instruct. The results, presented in Table 6, show that our method outperforms NTK-Aware Scaled RoPE in terms of accuracy and achieves lower latency, indicating the potential of our approach to scale effectively with larger models.
Method En.Sum En.QA R.KV
Acc. (%) Time (s) Acc. (%) Time (s) Acc. (%) Time (s)
NTK (SPDA) 23.49 199.52 28.77 145.69 50.00 111.98
TokenSelect 25.07 114.24 29.91 71.98 88.12 63.27
Table 6: Performance and latency comparison on Qwen2-72B-Instruct with tensor parallelism size: 4.
Appendix C Formal Statement and Proof of Lemma
@@
$\min_{1\leq i\leq k}{\hat{\mathbf{q}}_{2}\mathbf{k}_{i}}$ $-\max_{j>k}{\hat{\mathbf{q}}_{2}\mathbf{k}_{j}}$
$\geq\hat{\mathbf{q}}_{1}\mathbf{k}_{p}\cos{\theta}-\|\mathbf{k}\|_{\text{max}}\sin{\theta}$
$\quad\quad\quad\quad\quad\;\;-(\hat{\mathbf{q}}_{1}\mathbf{k}_{p+1}\cos{\theta}$
$\quad\quad\quad\quad\quad\;\;+\|\mathbf{k}\|_{\text{max}}\sin{\theta})$
@@
In order to have Eqn. (8) >0>0, we require
@@
$\Rightarrow\$$\frac{1-\cos^{2}{\theta}}{\cos^{2}{\theta}}<\Big(\frac{\hat{\eta}}{2\|\mathbf{k}\|_{\text{max}}}\Big)^{2},$
@@
Appendix D Overview of LLMs Inference
Nowadays, mainstream LLMs are primarily based on the Decoder-only Transformer architecture. Each transformer layer includes a multi-head attention (MHA) and a feed-forward networks (FFN). The inference process of LLMs can be divided into two stages: the prefill stage and the decode stage.
$\mathbf{Q}_{\text{prefill}},\mathbf{K}_{\text{prefill}},\mathbf{V}_{\text{prefill}}]=\mathbf{X}_{\text{prefill}}\cdot\left[\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}\right],$ (9)
$\mathbf{O}_{\text{prefill}}=\text{softmax}\left(\frac{\mathbf{Q}_{\text{prefill}}\cdot{\mathbf{K}_{\text{prefill}}}^{\top}}{\sqrt{d}}\right)\cdot\mathbf{V}_{\text{prefill}},\vskip-6.0pt$ (10)
where $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}$ are linear projections, $\cdot$] represents tensor concatenation operation, and Eq.(10) is also known as Scaled Dot Product Attention (SDPA). After these computation, $\mathbf{K}_{\text{prefill}}$ and $\mathbf{V}_{\text{prefill}}$ are stored as the KV Cache for current layer $\mathbf{K}_{\text{cache}}$ and $\mathbf{V}_{\text{cache}}$, and $\mathbf{O}_{\text{prefill}}$ is used for subsequent calculations.
Method En.Sum En.QA En.MC Math.F R.PK R.Num R.KV Avg.
H2O 2.8 0.7 0.0 6.0 2.5 2.4 0.0 2.1
InfLLM 24.3 19.5 43.7 23.7 100.0 99.0 5.0 45.0
TokenSelect 26.9 21.3 45.8 28.2 100.0 97.2 48.4 52.5
Table 7: Performance comparison with H2O Zhang et al. (2024b) on Llama-3-8B-Instruct, baseline performance is referenced from Xiao et al. (2024a).
Appendix E Comparison with Token Eviction-based Methods (e.g., H2O)
Token eviction–based methods Zhang et al. (2024b); Oren et al. (2024), led by H2Zhang et al. (2024b), have pioneered the field of long-context inference, achieving early state-of-the-art performance. Although both our method and H2O employ token‐level criticality estimation, they fall under two entirely different taxonomies. As discussed in Sec. 2 and Sec. 3, H2O is a query‐independent KV cache selection method, which suffers from three main drawbacks:
  1. 1.
    Lack of dynamism: Its importance scoring relies on attention scores from previous queries and keys. Consequently, KV pairs that are crucial for the current query may have been discarded earlier—a phenomenon also confirmed by QUEST Tang et al. (2024). Fig. 1 and 2 of QUEST provide an intuitive illustration of the differences between query‐based methods (e.g., our TokenSelect) and H2O. Notably, TokenSelect leverages a dynamic selection strategy, enabling state‐of‐the‐art performance with a minimal token budget.
  2. 2.
    Inability to extend sequence length: Since H2O depends on the model’s original attention mechanism, it cannot extend the effective context length. In contrast, our approach can easily extend a model with an original maximum length of 4K–32K tokens to an effective length exceeding 1M tokens.
  3. 3.
    Inefficient implementation: H2O evaluates token importance based on attention scores, making it incompatible with efficient kernels such as FlashAttention Dao (2024). This limitation restricts its scalability. Our method, however, is designed for broad compatibility and is fully transparent to large‐scale inference acceleration infrastructures, including paged attention, tensor parallelism, and prefix caching, making it ready for large‐scale online serving.
To further demonstrate the superiority of TokenSelect, we present experimental results in Table 7. These results corroborate the findings of previous studies Tang et al. (2024); Xiao et al. (2024a), showing that query‐independent methods are inferior to query‐based approaches.
Appendix F Detailed Descriptions on Baselines
@@
  • StreamingLLM Xiao et al. (2024b): The state-of-the-art method for long-context inference with predefined sparse patterns. Similar approaches include LM-Infinite Han et al. (2024).
@@
  • MInference Jiang et al. (2024): The state-of-the-art method for long-context prefilling acceleration, utilizing three sparse patterns including block-level sparse attention.
  • SnapKV (Li et al., 2024): A fine-tuning-free approach that efficiently compresses KV caches by selecting clustered important KV positions for each attention head.
  • InfiniGen Lee et al. (2024): A KV cache management framework that reduces memory overhead in offloading-based LLM inference by prefetching only essential KV cache entries through selective token rehearsal.
  • QUEST Tang et al. (2024): A query-aware KV cache management algorithm by selecting critical KV cache based on the query-aware sparsity at page granularity.
  • RetrievalAttention Liu et al. (2024a): The state-of-the-art method leveraging approximate nearest neighbor search on CPU memory and an attention-aware vector search algorithm to address distribution mismatches.
Appendix G More Information on Datasets
@@
Appendix H Comparison on Prefill Latency
@@
Length
FlashAttention-2
MInference
TokenSelect
@@
Appendix I Detailed Performance Comparisons Under Different Cache Threshold $\theta$
Table 9 presents the performance sensitivity to the threshold $\theta$ of the Selection Cache across various tasks. The results indicate that although $\theta$-sensitivity varies across different task types, most tasks exhibit only slight performance degradation as $\theta$ decreases. This suggests potential for further accelerating TokenSelect’s decode stage by reducing $\theta$ in the vast majority of cases. It is worth noting, however, that more challenging retrieval tasks—such as R.KV—show noticeable performance degradation as $\theta$ decreases, indicating higher dynamicity requirements for token selection in these tasks.
$\theta$ En.Sum En.QA En.MC En.Dia Code.D Math.F R.PK R.Num R.KV Avg.
0.5 20.99 17.83 54.31 7.50 30.20 21.14 100.00 96.10 0.20 38.69
0.6 21.21 18.08 54.31 7.50 30.20 21.36 100.00 96.78 0.20 38.84
0.7 20.73 18.08 54.31 7.50 30.46 21.36 100.00 98.98 4.40 39.53
0.8 21.47 17.85 54.31 7.50 30.20 21.58 100.00 100.00 24.60 41.94
0.85 22.39 18.15 54.31 7.50 30.20 21.79 100.00 100.00 68.20 46.94
0.9 22.62 18.86 54.31 7.50 30.20 21.71 100.00 100.00 86.60 49.08
0.95 22.46 18.54 54.31 7.50 30.56 21.77 100.00 100.00 86.20 49.05
1.0 22.66 18.68 54.31 7.50 30.51 21.78 100.00 100.00 86.84 49.15
Table 9: Performance using different selection cache similarity thresholds using Qwen2-7B-Instruct.
Appendix J Experimental Results on LongBench
@@ Appendix I Detailed Performance Comparisons Under Different Cache Threshold $\theta$
Appendix K Use of AI Assistants
In this paper, AI Assistants were used for literature retrieval and grammar checking.